OpenStudy (anonymous):

2/5(x+2)+1/4=1/2(5-3x). Can someone please explain and show steps on how to solve.

OpenStudy (anonymous):

PLEASE GIVE BEST RESPONSE. IT HELPS. _______________ P arentheses E xponents M ultiplication and D ivision A ddition and S ubtraction 2/5(x + 2) + 1/4 = 1/2(5 - 3x) __________________________________________________________ First lets do the parentheses using the distributive property. __________________________________________________________ 0.4x + 4/5 + 1/4 = 2.5 - 1.5x ___________________________________________________________ No Multiplication or Division so we go onto addition and subtraction, but first we need to make 4/5 and 1/4 have common denominators, which is 20. ____________________________________________________________ 0.4x +16/20 + 5/20 = 2.5 - 1.5x 0.4x + 21/20 = 2.5 - 1.5x 0.4x + 1, 1/20 = 2.5 -1.5x _____________________________________________________________ Now we will add 1.5x to both sides of the equal sign. _____________________________________________________________ 0.4x + 1, 1/20 = 2.5 -1.5x +1.5x _____________________________________________________________ This gets rid of the variable on one side (to make things simpler i will also turn 1 1/20 into a decimal. _____________________________________________________________ 1.9x + 1.05 = 2.5 _____________________________________________________________ We will now subtract 1.05 from both sides _____________________________________________________________ 1.9x + 1.05 - 1.05 = 2.5 - 1.05 _____________________________________________________________ 1.9x = 1.45 _____________________________________________________________ Now we will divide 1.9 by both sides and your done :) _____________________________________________________________ 1.9x/1.9 = 1.45/1.9 x = 0.76

OpenStudy (anonymous):

awesome. Thank you.